OpenStudy (anonymous):

a library expansion began in 2002, was expected to cost $127million. By 2006 library officials estimate the cost would be $39millon over budget. By what percent did the cost exceed the initial cost? HELLLPPP

OpenStudy (anonymous):

If y is a total, and x is some part of that total, then the percentage is 100x/y.

OpenStudy (anonymous):

WHAT?!

OpenStudy (anonymous):

If you have a total, say 150, then 60 is going to be what percentage of that? 60/150 = 0.4*100 = 40%. I'm trying to explain the concept, as it's a simple one and something you definitely need to know.

OpenStudy (anonymous):

dont understand i need help solving this problem

OpenStudy (anonymous):

the cost will be 39 million over budget, so the new cost is 127 + 39 = 166 million. There is a formula for percent increase.. new value - old value / old value and then multiply that by 100. Lets set it up.. 166 (new value) - 127 (old value) / 127 (old value) 39/127 = .307 .307 x 100 = 30.7% which rounds to 31% Do you understand ?